OpenStudy (anonymous):

An acorn falls from a tree. How fast is the acorn moving after 1.2 seconds?

OpenStudy (perl):

does it say to use meters per second for speed units, or ft/sec

OpenStudy (perl):

v = 9.8*t in m/s v = 32 * t in ft/s

OpenStudy (perl):

@AJ01 the initial velocity is zero , y = vo*t - 1/2 *at^2 , vo is initial velocity

OpenStudy (perl):

The general equation is y =-1/2 * a t^2 + vo*t + yo vo = initial velocity (here zero) yo = initial height (of the tree) y =- 1/2 * at^2 + 0*t + yo , y = -1/2 * at^2 + yo

OpenStudy (anonymous):

Yes its m/s.

OpenStudy (perl):

v= dy/dt = -1/2 *2* a * t = -a*t

OpenStudy (anonymous):

yes... you are right

OpenStudy (perl):

@janello do you see how to solve it? plug in 1.2 for t in v = 9.8 * t

OpenStudy (anonymous):

So 11.8? @Perl